# Periodicity and decidability of translational tilings by rational polygonal sets

### v1

@@ 1.1. Polygonal tiles
A polygonal set is a bounded open set $\Omega\subset\mathbb{R}^{2}$, whose boundary is a finite union of segments. Note that a polygonal set can certainly be disconnected and its connected components are not necessarily simply-connected. We denote the vertices of $\Omega$ by $V(\Omega)$, and the edges od $\Omega$ by $E(\Omega)$. We say that a polygonal set $\Omega$ is rational if the set $V(\Omega)-v$ is contained in $\mathbb{Q}^{2}$, for any $v\in V(\Omega)$.
Figure 1.1. Example of a polygonal set. The set $\Omega$ is shown in grey, the set $E(\Omega)$ consists of the 15 segments in black whose union is $\partial\Omega$, and $V(\Omega)$ is the 15 endpoints of those segments.
@@ 1.2. Outline of our argument
    Observe that by combining (2) and [GT20, Theorem 1.4], we obtain that any integer tiling in $\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ is weakly periodic. The main obstacle to transferring Theorem 1.2 from its integer counterpart is the possible sliding components obstructing periodicity. In Section 4, we address this obstacle by analysing the connection between the sliding components of any set in $\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ and the singly-periodic components of sets in $\operatorname{Tile}(F;N^{-1}\mathbb{Z}^{2})$, we conclude Theorem 1.2.
@@ 1.3. Notation
For a set $\Sigma$ in $\mathbb{R}^{d}$, we denote the closure of $\Sigma$ by $\overline{\Sigma}$.
@@ 2. From a tiling to an integer tiling
Let $\Omega$ be an integer tile. For $t,t^{\prime}$ in $\mathbb{R}^{2}$ we say that $t\approx_{\Omega}t^{\prime}$ if $(V(\Omega)+t)\cap(V(\Omega)+t^{\prime})$ is not empty. If $T$ is a tiling of $\mathbb{R}^{2}$ by $\Omega$, the relationship $\approx_{\Omega}$ induces (by transitive closure) an equivalence relationship in $T$. We denote this equivalence relationship by $t\sim_{\Omega,T}t^{\prime}$.
@@ 2.1. Reduction
The following lemma shows that to prove Lemma 2.2 it suffices to show that every polygonal tile can be repaired to a tile with a unique vertex sharing component. Thus, this will be our goal in the rest of the section.
@@
Let $\Omega$ be an integer tile, so that $T$ is a tiling of $\mathbb{R}^{2}$ by $\Omega$ containing 00. If $t\sim_{\Omega,T}t^{\prime}$ for any elements $t,t^{\prime}$ of $T$, then $T\subset\mathbb{Z}^{2}$.
@@
If $\Omega$ is an integer tile $t\approx_{\Omega}t^{\prime}$ implies that $t-t^{\prime}\in\mathbb{Z}^{2}$. Let $t\in T$. By assumption, since $0\in T$, we have $0\sim_{\Omega,T}t$. This means that there is a path $0=t_{1},t_{2},\dots,t_{n}=t$ with $t_{j}\in T$ and $t_{i}\approx_{\Omega}t_{i+1}$ for $i=1,\dots,n-1$. Therefore,
@@ 2.2. Sliding
by $\sim_{\Omega,T}$. Then there exists a unit vector $\vec{v}\in\mathbb{R}^{2}$ such that each set $W_{i}\coloneqq\Omega+T_{i}$ is $\vec{v}\mathbb{R}-$invariant.
@@
See also [GGRT23, Theorem 1.7] for a slightly related decomposition into “rational componenets” of tilings of a torus.
@@
We first show that if $x_{0}\in\partial(\Omega+S)$ is a proper vertex of $\Omega+S$, then $S$ is not $\sim_{\Omega,T}$-closed, i.e., there exist $t\in T\setminus S$ and $s\in S$ such that $s\approx_{\Omega}t$.
@@
and $T^{\prime}$ has a single $\sim_{\Omega,T}$-equivalence class.
@@
which ensures that an element of $T_{i}+s_{i}\vec{v}$ is $\sim_{\Omega,T}$-connected to an element of $T_{j}+s_{j}\vec{v}$, while preserving the $\sim_{\Omega,T}$-connectedness of $T_{i}$ and $T_{j}$. ∎
@@ 4.1. Earthquake decomposition
Let $F\subset N^{-1}\mathbb{Z}^{2}$ be finite. Suppose $T\in\operatorname{Tile}(F,N^{-1}\mathbb{Z}^{2})$, and $\vec{v}\in N^{-1}\mathbb{Z}^{2}$ then we define the $\vec{v}$-discrete earthquake decomposition of $T$ by $F$ as a decomposition
@@
Let $F\subset N^{-1}\mathbb{Z}^{2}$ be a tile of $N^{-1}\mathbb{Z}^{2}$, $T\in\operatorname{Tile}(F;N^{-1}\mathbb{Z}^{2})$ and $\operatorname{Plates}(T,\vec{v})$ be the $\vec{v}$-earthquake decomposition of $T$. Then there are pairwise incommensurable vectors $h_{1},\dots,h_{m}\in\mathbb{Z}^{2}$ such that for each $P\in\operatorname{Plates}(T,\vec{v})$ there is a decomposition $P=\bigsqcup_{j=1}^{m}P_{j}$ such that $P_{j}$ is $\langle h_{j}\rangle$-periodic for every $1\leqslant j\leqslant m$.
@@
Let $T^{\prime}\in\operatorname{Tile}(\tilde{\Omega};\mathbb{R}^{2})$ be the tiling that arises from merging the $\sim_{\tilde{\Omega},T}$ components of $T$ by sliding, as described in Lemma 2.7. Then, by construction, there are coefficients $s_{i}\in\mathbb{R}$ such that

### v2

@@ 1.1. Polygonal tiles
A polygonal set is a bounded open set $\Omega\subset\mathbb{R}^{2}$, whose boundary is a finite union of segments. Note that a polygonal set can certainly be disconnected and its connected components are not necessarily simply-connected. We denote the vertices of $\Omega$ by $V(\Omega)$, and the edges of $\Omega$ by $E(\Omega)$. See Figure 1.1 for illustration.
Figure 1.1. Example of a polygonal set $\Omega$. The set $\Omega$ is shown in gray; it has three connected components. The set $E(\Omega)$ consists of the $22$ segments in black whose union is the boundary of $\Omega$, $\partial\Omega$, and $V(\Omega)$ is the $21$ endpoints of those segments.
We say that a polygonal set $\Omega$ is rational if the set $V(\Omega)-v$ is contained in $\mathbb{Q}^{2}$, for any $v\in V(\Omega)$.
@@ 1.2. Outline of our arguments
    While the proof of Theorem 1.1 uses off-the-shelf results from [B20, GT20], the proof of Theorem 1.2 is more delicate and requires novel refinement of the technology introduced in [GT20]. Indeed, by combining (2) and [GT20, Theorem 1.4], we merely obtain that any integer tiling in $\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ is weakly periodic. The main obstacle to transferring Theorem 1.2 from its integer counterpart is the possible sliding components obstructing periodicity. In Section 4, we address this obstacle by analysing the connection between the sliding components of any set in $\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ and the singly-periodic components of sets in $\operatorname{Tile}(F;N^{-1}\mathbb{Z}^{2})$, and conclude Theorem 1.2.
@@ 1.3. Notation
For $\Sigma\subset\mathbb{R}^{d}$, we denote the closure of $\Sigma$ by $\overline{\Sigma}$ and the boundary of $\Sigma$ by $\partial\Sigma$.
@@ 1.4. Acknowledgements
We are grateful to Terence Tao and the anonymous referee for their helpful suggestions, which improved the exposition of this paper.
@@ 2. From a tiling to an integer tiling
Let $\Omega$ be an integer tile. For $t,t^{\prime}$ in $\mathbb{R}^{2}$ we say that $t\approx_{\Omega}t^{\prime}$ if $(V(\Omega)+t)\cap(V(\Omega)+t^{\prime})$ is not empty. If $T$ is a tiling of $\mathbb{R}^{2}$ by $\Omega$, the relationship $\approx_{\Omega}$ induces (by transitive closure) an equivalence relationship in $T$. We denote this equivalence relationship by $t\sim_{\Omega}t^{\prime}$.
@@ 2.1. Reduction
The following lemma shows that to prove Lemma 2.2 it suffices to show that every polygonal tile can be repaired to a tile with a unique vertex sharing component.
@@
Let $\Omega$ be an integer tile, so that $T$ is a tiling of $\mathbb{R}^{2}$ by $\Omega$ containing 00. If $t\sim_{\Omega}t^{\prime}$ for any elements $t,t^{\prime}$ of $T$, then $T\subset\mathbb{Z}^{2}$.
@@
If $\Omega$ is an integer tile $t\approx_{\Omega}t^{\prime}$ implies that $t-t^{\prime}\in\mathbb{Z}^{2}$. Let $t\in T$. By assumption, since $0\in T$, we have $0\sim_{\Omega}t$. This means that there is a path $0=t_{1},t_{2},\dots,t_{n}=t$ with $t_{j}\in T$ and $t_{i}\approx_{\Omega}t_{i+1}$ for $i=1,\dots,n-1$. Therefore,
@@ 2.2. Sliding
Lemma 2.4 reduces the proof of Lemma 2.2 to showing, for any integer tile $\Omega\subset\mathbb{R}^{2}$, the existence of a tiling $T\in\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ that is $\sim_{\Omega}$-connected, i.e., has a single $\sim_{\Omega}$-equivalence class. This will be achieved in this subsection, by developing a sliding machinery, which allows one to shift the $\sim_{\Omega}$-components of a given tiling $T\in\operatorname{Tile}(\Omega;\mathbb{R}^{2})$, while preserving the tiling, to eventually merge all of the components into a tiling $T^{\prime}\in\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ with a single $\sim_{\Omega}$-equivalence class (see Figure 2.2 for an illustration).
This is done by first analysing the structure of the $\sim_{\Omega}$-tiling-components, and showing that they are invariant under sliding in some direction $\vec{v}\in\mathbb{R}^{2}$, i.e., $\Omega+T_{i}+\delta\vec{v}=\Omega+T_{i}$, for every $\delta\in\mathbb{R}$ and $T_{i}\in T/\sim_{\Omega}$ (see Figure 2.1 for an illustration).
@@
by $\sim_{\Omega}$. Then there exists a unit vector $\vec{v}\in\mathbb{R}^{2}$ such that each set $W_{i}\coloneqq\Omega+T_{i}$ is $\vec{v}\mathbb{R}-$invariant.
Figure 2.1. In blue, green and orange, three sliding components, $W_{1},W_{2},W_{3}$, of a tiling, with a sliding direction $v$.
@@
See also [GGRT23, Theorem 1.7] for a related decomposition into “rational componenets” of tilings of a torus.
@@
We first show that if $x_{0}\in\partial(\Omega+S)$ is a proper vertex of $\Omega+S$, then $S$ is not $\sim_{\Omega}$-closed, i.e., there exist $t\in T\setminus S$ and $s\in S$ such that $s\approx_{\Omega}t$.
@@
Using Lemma 2.5, one can now slide the $\sim_{\Omega}$-components of a given tiling $T\in\operatorname{Tile}(\Omega;\mathbb{R}^{2})$ to merge them into a single $\sim_{\Omega}$-component of another tiling $T^{\prime}\in Tile(\Omega;\mathbb{R}^{2})$.
@@
and $T^{\prime}$ has a single $\sim_{\Omega}$-equivalence class.
Figure 2.2. The $\sim_{\Omega}$-classes from Figure 2.1 before sliding (left) and after sliding (right).
@@
which ensures that an element of $T_{i}+s_{i}\vec{v}$ is $\sim_{\Omega}$-connected to an element of $T_{j}+s_{j}\vec{v}$, while preserving the $\sim_{\Omega}$-connectedness of $T_{i}$ and $T_{j}$. ∎
@@ 3.2. Discretization
In this subsection, we encode any rational polygonal set $\Omega$ as a finite set in a two-dimensional lattice, in a way that preserves the tiling properties of $\Omega$. This will enable us to utilize the proof of the two-dimensional discrete periodic tiling conjecture [B20, GT20] to obtain Theorem 1.1 and deduce Corollary 1.1.
@@ 4.1. Earthquake decomposition
Let $F\subset N^{-1}\mathbb{Z}^{2}$ be finite. Suppose $T\in\operatorname{Tile}(F;N^{-1}\mathbb{Z}^{2})$, and $\vec{v}\in N^{-1}\mathbb{Z}^{2}$ then we define the $\vec{v}$-discrete earthquake decomposition of $T$ by $F$ as a decomposition
@@
Let $F\subset N^{-1}\mathbb{Z}^{2}$ be a tile of $N^{-1}\mathbb{Z}^{2}$, $T\in\operatorname{Tile}(F;N^{-1}\mathbb{Z}^{2})$ and $\operatorname{Plates}(T,\vec{v})$ be the $\vec{v}$-earthquake decomposition of $T$. Then there are pairwise incommensurable vectors $h_{1},\dots,h_{m}\in\mathbb{Z}^{2}\setminus\{0\}$ such that for each $P\in\operatorname{Plates}(T,\vec{v})$ there is a decomposition $P=\bigsqcup_{j=1}^{m}P_{j}$ such that $P_{j}$ is $\langle h_{j}\rangle$-periodic for every $1\leqslant j\leqslant m$.
@@
Let $T^{\prime}\in\operatorname{Tile}(\tilde{\Omega};\mathbb{R}^{2})$ be the tiling that arises from merging the $\sim_{\tilde{\Omega}}$ components of $T$ by sliding, as described in Lemma 2.7. Then, by construction, there are coefficients $s_{i}\in\mathbb{R}$ such that